OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). My answer is -2 Am I right?

OpenStudy (mathstudent55):

(f+g)(-1) = f(-1) + g(-1) = 2(-1) + 1 - 2(-1) - 3 = -2 You are correct